The director of a local grad school program wants your help to calculate the break-even number of students for a 7 quarter long program. Here is the Data

Total Tuition per student = $60,000

Variable Cost per student = $8,000

Administrative Overhead = $1,000,000

Faculty Cost = $25,000 per quarter

A) What is Qee ?

B) the faculty wants a 20% raise. If the Director wants Qee to remain the same , and she can only change V,

what must the new value for V be to give the faculty the raise ?
Answer #1

a) Contribution per unit = Tuition fee - Variable cost = 60,000 -8,000 = 52,000

Fixed costs = 1,000,000 + (25,000*7) = 1,175,000

Break even point = 1,175,000 / 52,000 = 23 students ( approx.)

b) Increased fixed costs = 1,000,000 + ( 25,000 *120%*7) = 1,210,000

Contribution required if she can only change v = 1,210,000 / 23 = 52608.69565

Variable cost = 60,000 - 52608.69565 = 7391.3